PLEASE I NEED HELP ON THIS PROBLEMS: (3X^2-6)^2=
X^4-16=
(X+7)(X+9)=
X^2+14X+33=

Answer by stanbon(75887)

(3X^2-6)^2= 9x^4-36x^2+36
X^4-16= =(x^2+4)(x+2)(x-2)
(X+7)(X+9)= (x^2+16x+63)
X^2+14X+33= (x+11)(x+3)
-------------
Cheers,
Stan H.